# Hulk: A Universal Knowledge Translator for Human-Centric Tasks

### v4

@@ V-A1 Training Datasets
Specialist DETR [93] ViT-B 75.9 73.2 74.4
PEDR [94] ViT-B 91.6 43.7 83.3
DDETR [95] ViT-B 91.5 43.7 83.1
Sparse-RCNN [96] ViT-B 91.3 44.8 81.3
Iter-DDETR [97] ViT-B 92.1 41.5 84.0
Iter-Sparse-RCNN [97] ViT-B 92.5 42.6 83.3
@@
V-A2 Evaluation Datasets
V-B Implementation Details
Following UniHCP [2] and PATH [30], we use the standard ViT-Base [98] as the default backbone and initialize it with unsupervised MAE pretraining [99]. Unless otherwise specified, the input resolution is consistent within each task, i.e., 256$\times$192 resolution for 2D pose estimation and attribute recognition, 480$\times$480 resolution for human parsing, 224$\times$224 resolution for 3D pose estimation and mesh recovery, 17 joints×$\times$175 frames for skeleton-based action recognition, 384$\times$384 resolution for image caption, and a maximum height/width of 1120 for pedestrian detection. Due to the high GPU memory demand of detection tasks, we resize the maximum height/width of the input images from the commonly used 1333 to 1120, which has a certain impact on the performance of detection tasks but does not affect the superiority of Hulk over other methods.
Considering the computational and communication efficiency during training, we adopt the practices in UniHCP to run one specific task on each GPU with gradient checkpointing [100]. We use the same Adafactor optimizer as that in UniHCP with $\beta_{1}=0.9$, $\beta_{2}$ clipped at 0.999 and disable the parameter scaling. We adopt a 60000-iteration schedule with a linear warm-up for the first 2000 iterations. We set the base learning rate to $10^{-3}$ with a cosine decay scheduler. For ViT-B backbone, we use a drop-path-rate of 0.2, a layer-wise learning rate decay of 0.75, and set weight decay to 0.05. We also train the ViT-L with a drop-path-rate of 0.5, a layer-wise learning rate decay of 0.85, and a weight decay of 0.1, following the recommendation setup in ViTDET [101]. To better utilize the knowledge in every dataset, we do not restrict that all datasets should take the same epoch during training and provide detailed joint training setup in the supplemental material. The whole training takes 70, 144 hours in total on 80 NVIDIA A100 GPUS for ViT-B, ViT-L, respectively.
TABLE V: 2D Pose estimation evaluation on COCO [72] val set and on AIC [73] test set with mAP. We compare our method with other SOTA methods, both specialists and generalists, with an input resolution of 256$\times$192. †indicates the implementation. ‡indicates that multiple datasets are used.
@@ V-B Implementation Details
Specialist HRNet [102] HRNet-W32 74.4 -
HRNet [102] HRNet-W48 75.1 -
@@
TABLE VI: Human parsing evaluation on Human3.6m [71], LIP [75] and CIHP [74] with mIoU. We compare our method with other SOTA methods, including specialists, generalists, and pretrainings, with an input resolution of 480$\times$480. †indicated the re-implementation.
@@
Specialist SNT [106] ResNet-101 - 54.73 60.87
PCNet [107] ResNet-101 - 57.03 61.05
SCHP [80] ResNet-101 - 59.36 -
@@
V-C Comparison with States-of-the-art Methods
Extensive evaluations are conducted to demonstrate the capabilities of our Hulk across diverse human-centric tasks. Table IV-IX show the experimental results of Hulk on eight different human-centric tasks, i.e., pedestrian detection, 2D pose estimation, human parsing, pedestrian attribute recognition, 3D human pose and mesh recovery, skeleton-based action recognition, and pedestrian image captain.
Following the practice in UniHCP, we report two kinds of evaluation results of our Hulk: (1) direct evaluation, where the pre-trained models are directly used for evaluation on the target dataset, and (2) finetuning, where the pretrained Hulk is first finetuned with the train split of the target dataset and then evaluated.
V-C1 2D Vision Tasks
[…]
@@ V-C Comparison with States-of-the-art Methods
2D pose estimation. To evaluate the performance of Hulk on 2D pose estimation task, we conduct evaluations on both the COCO and AIC datasets. The results are listed in Table V. Hulk model improves previous SOTA by considerable margins on AIC, i.e., +0.6% AP and +0.8% on ViT-B and ViT-L, respectively. The performance gain demonstrates the good scalability of Hulk. Hulk achieves on-par performance compared with ViTPose++ [105] on COCO. We attribute the relatively small performance gain on COCO to the distribution gap between the training subset and evaluation subset, where models are trained with on ground-truth boxes but evaluated on estimated boxes with a detector obtaining human AP of 56 on COCO valset [113].
TABLE VIII: Skeleton-based Action Recognition evaluation on NTU60-XSub [114] with accuracy results reported.
Method Backbone Accuracy
Specialist ST-GCN [115] GCN 81.5
AS-GCN [116] GCN 86.8
AGCN [117] GCN 88.5
Shift-GCN [118] GCN 90.7
CrosSCLR [119] GCN 86.2
MCC [120] GCN 89.7
SCC [121] GCN 88.0
UNIK [122] GCN 86.8
CTR-GCN [123] GCN 92.4
MS-G3D [124] GCN 91.5
IndRNN [125] RNN 81.8
MTCNN [126] CNN 81.1
HCN [127] CNN 86.5
PoseConv3D [128] CNN 93.1
ST-TR [129] Transformer 89.9
DSTA-Net [130] Transformer 91.5
STTFormer [131] Transformer 89.9
Pretraining MotionBERT [1] DSTformer 93.0
Generalist Hulk ViT-B 93.8
Hulk-FT ViT-B 94.0
Hulk ViT-L 94.1
Hulk-FT ViT-L 94.3
Human parsing. We evaluate the performance of Hulk on the human parsing task on Human3.6M, LIP, and CIHP datasets. As shown in Table VI, when using ViT-B, Hulk surpasses SOTA methods by +2.36% mIoU, +1.46% mIoU on Human3.6M and CIHP, respectively. On LIP, although Hulk achieves a relatively small performance gain with ViT-B, Hulk showcases considerable performance with ViT-B, i.e., 66.02% mIoU, demonstrating excellent scalability of Hulk.
Pedestrian attribute recognition. We evaluate the performance of our Hulk on PA-100K and RAPv2 datasets to demonstrate its effectiveness on the pedestrian attribute recognition task. Results are listed in Table VII. As can be seen, on ViT-B, Hulk outperforms previous methods by obtaining 87.85% mA and 85.26% mA on PA-100K and RAPv2, respectively. The performance of the large Hulk with ViT-B further increases to 88.97% mA and 85.86% mA, showing competitive results without adopting task-specific heads in human-centric pretraining models [30].
@@ V-C2 Skeleton-based Tasks
In the skeleton-based action recognition task, existing methods utilize three categories of backbones: (1) Graph Convolutional Networks (GCN), which are the most common architecture; (2) Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN), and (3) more recently, Transformer architectures. Our Hulk falls within the third category. To assess the effectiveness of Hulk, we report results on NTU60-XSub, a well-known dataset on skeleton-based action recognition. The results in Table VIII indicate that even without specific pretraining designs [1], Hulk obtains a remarkable accuracy of 93.8% on ViT-B, better than the SOTA PoseConv3D [128] method. After finetuning Hulk with ViT-L backbone, the performance on NTU60-XSub further increases to 94.3%. This improvement not only highlights the effectiveness of Hulk in the skeleton-based action recognition, but also demonstrates Hulk’s ability of tokenizing skeleton and image inputs into a modality-shared manifold space to extract generalized human-centric knowledge.
TABLE IX: Pedestrian image captain evaluation on CUHK-PEDES [90].
Model Backbone B@4 CIDEr
Specialist BLIP [132] ViT-B 32.9 103.8
BLIP-2 [54] ViT-B 32.8 105.8
Generalist Hulk ViT-B 31.1 91.4
Hulk-FT ViT-B 28.3 72.8
Hulk ViT-L 31.6 94.5
Hulk-FT ViT-L 30.5 82.8
V-C3 Vision-language Tasks
We also evaluate Hulk on the CUHK-PEDES dataset for the vision-language task, i.e., pedestrian image captain. For fair comparisons, we re-implemented the state-of-the-art BLIP [132] and BLIP-2 [54] methods by finetuning its pre-trained weights (on a 129M image-text dataset) on the CUHK-PEDES dataset. As shown in Table IX, Hulk achieves on-par performance with BLIP and BLIP-2 on the B@4 metric but lags behind in terms of CIDEr. This discrepancy in performance can be attributed to the extensive pretraining of BLIP and BLIP-2 on large-scale vision-language datasets, while our Hulk is trained solely on smaller human-centric vision-language datasets (about 12M). Furthermore, we observed that finetuning on the CUHK-PEDES dataset led to a decline in the model’s prediction accuracy, because better pedestrian captions can learn from diverse annotations in large-scale human-centric datasets.
V-C4 3D Vision Tasks
TABLE X: monocular 3D human pose and mesh recovery evaluation on 3DPW [70] and Human3.6M [71] among image-based methods.
@@ V-C3 Vision-language Tasks
Specialist HMR [134] ResNet-50 - 130.0 76.7 88.0 56.8
GraphCMR [135] ResNet-50 - - 70.2 - 50.1
SPIN [136] ResNet-50 116.4 96.9 59.2 62.5 41.1
I2LMeshNet [137] ResNet-50 - 93.2 57.7 55.7 41.1
PyMAF [138] ResNet-50 110.1 92.8 58.9 57.7 40.5
ROMP [139] ResNet-50 105.6 89.3 53.5 - -
ROMP [139] HRNet-W32 103.1 85.5 53.3 - -
PARE [140] ResNet-50 99.7 82.9 52.3 - -
METRO [39] ResNet-50 - - - 56.5 40.6
METRO [39] HRNet-W64 88.2 77.1 47.9 54.0 36.7
@@ V-D Ablation Study
To demonstrate the effectiveness of Hulk, we conduct several ablations on a smaller training set. The ablation set consists of 21 datasets with about 1.7M samples, covering all eight tasks. We provide the detailed dataset combination in the supplemental material. Without otherwise specified, we train Hulk for 20k iterations in all ablation variants. While results on different datasets within the same task are highly related, we showcase one dataset per task to represent the performance on the specific task in the main text (all results are provided in the supplemental material). As the metric in 3D pose estimation and mesh recovery computes the error, we use $100-error$ when we need to compute average performance among 8 tasks.
@@ V-D1 Weight sharing
As Hulk demonstrates promising performance in various human-centric tasks while sharing mostly all parameters, we explore the impact of incorporating more task-specific parameters into our learning framework. To assess the effectiveness of our weight-sharing approach, we ablate three weight-sharing variants: (a) Our Hulk model, all parameters in both the encoder and decoder are shared across all datasets, and parameters in the tokenizers and de-tokenizers are shared within the same input/output modality. (b) All parameters in both the encoder and decoder are shared across all datasets, but the parameters in the tokenizers and de-tokenizers are dataset-specific. (c) Only encoder parameters are shared. In Table XI, we observed that more parameter sharing enhances the overall performance of the model. Specifically, introducing decoder parameter sharing (from (b) to (c)) significantly benefits task performances by +12.1% on average. This shows that shared human semantic information in different human-related annotations is beneficial to human-centric tasks. A notable example is the skeleton-based action recognition task. When only the encoder parameters are shared, the model fails to converge because of the small size of the NTU60-XSub dataset, resulting in only $\textbf{1.7}\%$ accuracy. However, by sharing more model parameters, the general human-centric knowledge from other tasks can be substantially beneficial to better accuracy in this task.
TABLE XI: Comparison of different parameter-sharing schemes. $\mathcal{E}$, $\mathcal{D}$, $\mathcal{P}$, and $\mathcal{Q}$ denote encoder, decoder, tokenizer, and de-tokenizer, respectively. With more task-specific parameters, the average tasks performance of Hulk decline significantly.
@@
LIP AIC CrowdHuman RAPv2 CUHK-PEDES NTU60-Sub 3DPW 3DPW

### v5

@@ V-B Implementation Details
Considering the computational and communication efficiency during training, we adopt the practices in UniHCP to run one specific task on each GPU with gradient checkpointing [119]. We use the Adafactor optimizer with $\beta_{1}=0.9$, $\beta_{2}$ clipped at 0.999 and disable the parameter scaling. We adopt a 60000-iteration schedule with a linear warm-up for 2000 iterations. We set the base learning rate to $10^{-3}$ with a cosine decay scheduler. For ViT-B backbone, we use a drop-path-rate of 0.2, a layer-wise learning rate decay of 0.75, and set weight decay to 0.05. For ViT-L, we use a drop-path-rate of 0.5, a layer-wise learning rate decay of 0.85, and a weight decay of 0.1, following the recommendation setup in ViTDET [120]. To better utilize the knowledge in every dataset, we do not restrict all datasets to take the same epoch. As different task loss scales vary, to prevent model bias to a certain task, we monitor gradients on task-shared parameters, ensure similar gradients for all tasks, and ablate them for better performance. Joint training setups are detailed in Sec. VIII of the supplemental material. The whole training takes 70, 144 hours in total on 80 NVIDIA A100 GPUS for ViT-B, ViT-L, respectively.
V-C Comparison with States-of-the-art Methods
Extensive evaluations are conducted to demonstrate the capabilities of our Hulk across diverse human-centric tasks. Table XXVII-X show the experimental results of Hulk on eight different human-centric tasks, i.e., pedestrian detection, 2D pose estimation, human parsing, pedestrian attribute recognition, 3D human pose and mesh recovery, skeleton-based action recognition, and pedestrian image caption.
Following UniHCP, we report two kinds of evaluation results of Hulk: (1) direct evaluation, where the pre-trained models are directly used for evaluation on the target dataset, and (2) finetuning, where the pretrained Hulk is first finetuned with the train split of the target dataset and then evaluated.
V-C1 2D Vision Tasks
TABLE IV: Pedestrian detection evaluation on CrowdHuman [102] with mAP, MR-2 and JI. Following UniHCP [2], we report the direct eval and fine-tune (FT) results. †indicates using a smaller input resolution with a maximum height/width of 1120.
@@ V-C Comparison with States-of-the-art Methods
Specialist DETR [121] ResNet-50 75.9 73.2 74.4
CrowdDet [122] ResNet-50 90.7 41.4 82.3
V2F-Net [123] ResNet-50 91.0 42.3 -
PEDR [124] ResNet-50 91.6 43.7 83.3
DDETR [125] ResNet-50 91.5 43.7 83.1
Sparse-RCNN [126] ResNet-50 91.3 44.8 81.3
Iter-DDETR [47] ResNet-50 92.1 41.5 84.0
Iter-Sparse-RCNN [47] ResNet-50 92.5 42.6 83.3
@@
[…]
@@
Specialist HRNet [43] HRNet-W48 75.1 -
@@
TABLE VI: Human parsing evaluation (mIoU) on Human3.6m [96], LIP [62] and CIHP [99]. We compare Hulk with other SOTA methods, including specialists, generalists, and pretrainings, with an input resolution of $\times480$. †indicated re-implementation.
@@
Specialist SCHP [104] ResNet-101 - 59.36 -
@@
Pedestrian detection. In the pedestrian detection task, MR-2 is employed as a more sensitive and crucial metric in detection in the crowd scenes. This metric is particularly sensitive to False Positives (FP), with lower values indicating more desirable performance. We use an input image resolution of a maximum height/weight of 1120, while other methods take a large resolution of 1333 by default. As shown in Table  XXVII, Hulk showcases competitive performance with State-Of-The-ART (SOTA) using a smaller input resolution. Specifically, after finetuning, our Hulk achieves an improvement of -0.7% MR-2 on ViT-B and -1.2% MR-2 on ViT-L, respectively, compared to SOTA specialist methods, e.g., Iter-DDETR [47].
@@
2D pose estimation. To evaluate the performance of Hulk on 2D pose estimation task, we conduct evaluations on both the COCO and AIC datasets. The results are listed in Table XXVIII. Hulk improves previous SOTA by considerable margins on AIC, i.e., +0.6% and +0.8% AP on ViT-B and ViT-L, respectively. The performance gain demonstrates the good scalability of Hulk. Hulk achieves on-par performance compared with ViTPose++ [130] on COCO. We attribute the relatively small performance gain on COCO to the distribution gap between the training set and val set, where models are trained with on ground-truth boxes but evaluated on estimated boxes with a detector obtaining human AP of 56 on COCO val set [131].
TABLE VIII: Skeleton-based Action Recognition evaluation on NTU60-XSub [132] and FineGYM [112] with accuracy reported.
Method Backbone NTU60 GYM
Specialist ST-GCN [48] GCN 81.5 25.2
Shift-GCN [54] GCN 90.7 -
CTR-GCN [53] GCN 92.4 -
MS-G3D [133] GCN 91.5 92.0
PoseConv3D [134] CNN 93.1 92.4
SELFYNet [135] CNN - 87.7
ST-TR [55] Transformer 89.9 -
DSTA-Net [56] Transformer 91.5 -
STTFormer [57] Transformer 89.9 -
Pretraining MotionBERT [1] DSTformer 93.0 -
SkeletonMAE [136] GIN - 91.8
Generalist Hulk ViT-B 93.8 91.6
Hulk-FT ViT-B 94.0 92.2
Hulk ViT-L 94.1 92.3
Hulk-FT ViT-L 94.3 93.2
Human parsing.
We evaluate the performance of Hulk on the human parsing task on Human3.6M, LIP, and CIHP datasets. As shown in Table XXIX, when using ViT-B, Hulk surpasses SOTA methods by +2.36% mIoU, +1.46% mIoU on Human3.6M and CIHP, respectively. On LIP, although Hulk achieves a relatively small performance gain with ViT-B, Hulk showcases considerable performance with ViT-B, i.e., 66.02% mIoU, demonstrating excellent scalability of Hulk.
Pedestrian attribute recognition.
We evaluate the performance of Hulk on PA-100K and RAPv2 to demonstrate its effectiveness on the pedestrian attribute recognition task. As shown in Table XXX, on ViT-B, Hulk outperforms previous methods by obtaining 87.85% mA and 85.26% mA on PA-100K and RAPv2, respectively. The performance of the large Hulk with ViT-B further increases to 88.97% mA and 85.86% mA, showing competitive results without adopting task-specific heads in human-centric pretraining models [30].
@@ V-C2 Skeleton-based Tasks
In the skeleton-based action recognition task, existing methods utilize three categories of backbones: (1) Graph Convolutional Networks (GCN), which are the most common architecture; (2) Convolutional Neural Networks (CNN) and Recurrent Neural Networks (RNN), and (3) Transformer architectures. Our Hulk falls within the third category. To assess the effectiveness of Hulk, we report results on NTU60-XSub, a well-known dataset on skeleton-based action recognition. The results in Table XXXI indicate that even without specific pretraining designs [1], Hulk obtains a remarkable accuracy of 93.8% on ViT-B, better than the SOTA PoseConv3D [134] method. After finetuning Hulk with ViT-L backbone, the performance on NTU60-XSub further increases to 94.3%. This improvement not only highlights the effectiveness of Hulk in skeleton-based action recognition, but also demonstrates Hulk’s ability of tokenizing skeleton and image inputs into a modality-shared manifold space to extract generalized human-centric knowledge. Using the skeleton data in the FineGYM dataset as input, Hulk is also better than the state-of-the-art methods, achieving the accuracies of 92.2 and 93.2 on ViT-B and ViT-L, respectively.
TABLE IX: Monocular 3D human pose and mesh recovery evaluation on 3DPW [95] and Human3.6M [96] among image-based methods.
@@
Specialist METRO [39] HRNet-W64 88.2 77.1 47.9 54.0 36.7
@@
TABLE X: Pedestrian image caption evaluation on CUHK-PEDES [114]. †denotes fixed backbone. PD - Pretrain dataset. TM - whether has pretrained Text Models or not.
Model Backbone Params PD TM B@4 CIDEr
Specialist BLIP [141] ViT-B 446.0M 129M 32.9 103.8
BLIP-2 [79] ViT-G† 1.1B 129M 32.8 105.8
OFA-L [92] ResNet-101 472.0M 18M $\times$ 31.5 97.7
mPLUG-B [142] ViT-B 350.0M 14M 32.2 101.1
Generalist Hulk ViT-B 113.7M 12M $\times$ 31.1 91.4
Hulk-FT ViT-B 113.7M 12M $\times$ 28.3 72.8
Hulk ViT-L 336.0M 12M $\times$ 31.6 94.5
Hulk-FT ViT-L 336.0M 12M $\times$ 30.5 82.8
V-C3 Vision-language Tasks
We also evaluate Hulk on the CUHK-PEDES dataset for the vision-language task, i.e., pedestrian image caption. For fair comparisons, we re-implemented the state-of-the-art methods, i.e., OFA [92], mPLUG [142], BLIP [141] and BLIP-2 [79], by finetuning their pre-trained weights on the CUHK-PEDES dataset. As shown in Table X, Hulk achieves on-par performance with SOTA methods on the B@4 metric but lags behind in terms of CIDEr. This discrepancy in performance can be attributed to two factors: (1) the extensive pretraining of BLIP and BLIP-2 on large-scale vision-language datasets, while our Hulk is trained solely on smaller human-centric vision-language datasets (about 12M). (2) For unifying other human-centric tasks, Hulk does not incorporate a pretrained text encoder/decoder, a crucial module for generating coherent text outputs in modern caption models. Furthermore, we observed that finetuning on the CUHK-PEDES dataset led to a decline in the model’s prediction accuracy, because better pedestrian captions can learn from diverse annotations in large-scale human-centric datasets.
V-C4 3D Vision Tasks
@@ V-D Ablation Study
To demonstrate the effectiveness of Hulk, we conduct several ablations on a smaller training set. The ablation set consists of 21 datasets with about 1.7M samples, covering all eight tasks. We provide the detailed dataset combination in the supplemental material. Without otherwise specified, we train Hulk for 20k iterations in all ablation variants. While results on different datasets within the same task are highly related, we showcase one dataset per task to represent the performance on the specific task in the main text (all results are provided in the supplemental material). As the metric in 3D pose estimation and mesh recovery computes the error, we use 100error100-error when we need to compute average performance among 8 tasks.
@@ V-D1 Weight sharing
As Hulk demonstrates promising performance in various human-centric tasks while sharing mostly all parameters, we explore the impact of incorporating more task-specific parameters into our learning framework. To assess the effectiveness of our weight-sharing approach, we ablate three weight-sharing variants: (a) Our Hulk model, all parameters in both the encoder and decoder are shared across all datasets, and parameters in the tokenizers and de-tokenizers are shared within the same input/output modality. (b) All parameters in both the encoder and decoder are shared across all datasets, but the parameters in the tokenizers and de-tokenizers are dataset-specific. (c) Only encoder parameters are shared. In Table XI, we observed that more parameter sharing enhances the overall performance of the model. Specifically, introducing decoder parameter sharing (from (b) to (c)) significantly benefits average performances by +12.1%. This shows that shared human semantic information in different human-related annotations is beneficial to human-centric tasks. A notable example is the skeleton-based action recognition task. When only the encoder parameters are shared, the model fails to converge because of the small size of the NTU60-XSub dataset, resulting in only $\textbf{1.7}\%$ accuracy. However, by sharing more model parameters, the general human-centric knowledge from other tasks can be substantially beneficial to better accuracy in this task.
TABLE XI: Comparison of different parameter-sharing schemes. $\mathcal{E}$, $\mathcal{D}$, $\mathcal{P}$, and $\mathcal{Q}$ denote encoder, decoder, tokenizer, and de-tokenizer, respectively. With more task-specific parameters, the average tasks performance of Hulk declines significantly.
@@
LIP AIC CrowdHuman RAPv2 CUHK-PEDES NTU60-XSub 3DPW 3DPW
@@
TABLE XII: Ablation of collaboration and interference between tasks. In green and red are the gaps of at least $\pm 0.5$ points.
LIP AIC CrowdHuman RAPv2 CUHK-PEDES NTU60-XSub 3DPW 3DPW
mIoU AP AP mA B@4 Acc 100-MPVPE 100-MPJPE
baseline 56.8 25.5 81.8 77.9 28.0 93.2 -0.3 13.2 -
w/o Parsing - 25.2↓0.3 78.5↓3.3 77.3↓0.6 25.8↓2.2 93.2↑0.0 -2.9↓2.6 10.8↓2.4 -1.6
w/o 2D Pose 52.3↓4.5 - 65.0↓16.8 75.1↓2.8 26.6↓1.5 93.5↑0.3 -18.9↓18.6 -3.5↓16.7 -8.6
w/o Detection 57.6↑0.8 26.4↑0.9 - 77.6↓0.3 27.8↓0.2 93.1↓0.1 -4.8↓4.5 10.5↓2.7 -0.9
w/o Attribute 56.9↑0.1 25.6↑0.1 79.9↓1.9 - 28.2↑0.2 93.1↓0.1 -2.1↓1.8 13.2↓1.0 -0.6
w/o Caption 57.3↑0.5 25.6↑0.1 80.2↓1.6 77.8↓0.1 - 92.9↓0.3 -2.2↓1.9 11.2↓2.0 -0.8
w/o Skeleton 57.4↑0.6 25.7↑0.2 80.9↓0.9 77.4↓0.5 28.2↑0.2 - -4.8↓4.5 9.9↓3.3 -1.2
w/o Mesh&3D Pose 57.4↑0.6 25.5↑0.0 81.0↓0.8 78.0↑0.1 27.9↓0.1 93.1↓0.1 - - -0.1
TABLE XIII: Ablation of how generic tasks help smaller ones.
Task Detection Attribute Skeleton
CrowdHuman RAPv2 NTU60-XSub
AP mA Acc
baseline 52.6 75.4 93.1
baseline + 2D pose joint train 71.6 78.7 93.5
baseline + 3D pose joint train 57.8 76.5 93.2